OpenStudy (anonymous):

The average speed of an express train is 15mi/hr faster than the average speed of a local train. It takes the local train 2 hours longer than the express train to cover 360 miles. What is the average speed of the express train?

OpenStudy (kropot72):

Let speed of local train = x Therefore speed express = x + 15 Time for local train to do 360 miles = 360/x Time for express to do 360 miles =\[\frac{360}{x+15}\]Subtracting 2 from the time for the local train gives a time equal to the express: \[\frac{360}{x}-2=\frac{360}{x+15}\] This above simplifies to a quadratic as follows: \[2x ^{2}+30x-5400=0\] The next step is to solve for x

OpenStudy (kropot72):

The solution to the quadratic is x = 45 miles per hour for the speed of the local train. Therefore the speed of the express train is 45 +15 = 60 miles per hour.

OpenStudy (anonymous):

can you solve doing a chart?

OpenStudy (kropot72):

Not really. Can't you follow the algebraic method? Note: At the present time the equation editor on this web site is giving me problems and the working can no longer be followed. Maybe you have the same trouble

OpenStudy (anonymous):